OpenStudy (anonymous):

If each balloon is filled with carbon dioxide gas at 20ºC and 1 atmosphere, calculate the mass and the number of moles of carbon dioxide in each balloon at maximum inflation. Use the ideal gas law in your calculation

OpenStudy (anonymous):

Balloon: 1.) 15 2.) 16 3.) 19

OpenStudy (abb0t):

Use the ideal gas law to solve for moles, \(n\). See, you have: \(\sf \color{blue}{PV=nRT}\), if you rearrange that to solve for \(n\), you get \(\Rightarrow\) \(\sf \color{red}{\frac{PV}{RT} = n = moles}\) now, once you have moles, you can solve for mass by multiplying the moles by the molar mass, the units cancel out: \(\sf \color{violet}{\cancel{moles~CO_2} \times \frac{mass~CO_2}{\cancel{1~mole~CO_2}}}=grams~of~CO_2\)

OpenStudy (anonymous):

Thank you so much!!! @abb0t